Andrew has more money than Ben, if Andrew gave Ben £20, they would have the same amount, while if Ben gave Andrew £22, Andrew would have then have twice as much as Ben.
So, new working
A-20 = B+20
B-22 =2A +22
Solve for one, substitute new value into other equation. I "solved" equation 2, and got:
B=2A+44
substituting it into equation 1:
A-20=2A+44+20
A=2A+84
A=84
I ended up with a result of 84 (wrong)

Answer by Theo(13342)


the problem states:

Andrew has more money than Ben, if Andrew gave Ben £20, they would have the same amount, while if Ben gave Andrew £22, Andrew would have then have twice as much as Ben.

Andrew has more money than Ben leads to:

A > B.

if Andrew gave Ben 20, they would have the same amount leads to:

A - 20 = B + 20

if Ben gave Andrew 22, Andrew would have twice as much as Ben leads to:

A + 22 = 2 * (B - 22).

you have:

A - 20 = B + 20
A + 22 = 2 * (B - 22)

simplify to get:

A - 20 = B + 20
A + 22 = 2B - 44

get all the A's and B's on the same side of the equation and you get:

A - B = 40
A - 2B = -66

subtract the second equation from the first and you get:

B = 106

use either of the first two original equations to find the value of A.

i used the second one.

A + 22 = 2 * (B - 22)
when B = 106, this becomes A + 22 = 2 * (106 - 22).
subtract 22 from both side and simplify to get A = 146.

you have A = 146 and B = 106

evaluate both original equations to see if they hold true with these values for A and B.

A - 20 = B + 20 becomes 146 - 20 = B + 20 which becomes 126 = 146 which is true.

A + 22 = 2 * (B - 22) becomes 146 + 22 = 2 * (106 - 22) which becomes 168 = 2 * (84) which becomes 168 = 168 which is true.

the values for A and B are good.

your solution is that A = 146 and B = 106.